OpenStudy (anonymous):

. Jeremy is building a large deck for a community center. The deck is shaped as a rectangle. The width of the deck is 29 feet. The perimeter of the deck is to be at least 134 feet. a. Write an inequality that represents all possible values for the length of the deck. b. Find all possible values for the length of the deck.

OpenStudy (anonymous):

I just need help solving it, I don't want the answer just need help :P

OpenStudy (hoblos):

 let L represent length and W represent width perimeter = L +W +L +W = 2L + 2W the deck is to be at least 134 feet, thus perimeter is greater or equal 134

OpenStudy (hoblos):

can you solve it now ?

OpenStudy (anonymous):

im trying one sec.

OpenStudy (hoblos):

ok, tell me what you get to check it for you

OpenStudy (anonymous):

so 58W*X2?

OpenStudy (hoblos):

what's X ?

OpenStudy (mathmale):

"let L represent length and W represent width perimeter = L +W +L +W = 2L + 2W" Very nice. Note that the width of the deck is a fixed 29 feet. Substitute 29 ft. for W. @hoblos: "the deck is to be at least 134 feet, thus perimeter is greater or equal 134" doesn't make sense. Please reread the original question and fix this. @letsparty: Where did "so 58W*X2?" come from? Please explain your reasoning. Why not use W and L as your variables?

OpenStudy (anonymous):

I did 29+29 to get my width

OpenStudy (anonymous):

sorry 58W*L2

OpenStudy (hoblos):

ohh you're right @mathmale my bad i forgot 2 words ! so it is "the perimeter of the deck is to be at least 134 feet, thus perimeter is greater or equal 134"

OpenStudy (mathmale):

@letsparty: The width (W) of the deck is 29. Thus, the perimeter will be twice that, plus twice the length. Write that out in symbols, please.

OpenStudy (hoblos):

it is 2L + 2W @letsparty now replace W by 29 what do you get ?

OpenStudy (anonymous):

29+29+L+L=58*L*L

OpenStudy (mathmale):

Right start. But where does this multiplication ( * ) come from?

OpenStudy (anonymous):

so just 58+L+L

OpenStudy (hoblos):

yeah now it's correct

OpenStudy (mathmale):

Yes, or 58+2L. Please label this expression. What's it for?

OpenStudy (anonymous):

to find the perimeter of the deck

OpenStudy (mathmale):

Right. @hoblos, what next?

OpenStudy (hoblos):

so we find the expression of the perimeter, and we already figured out that it has to be greater or equal to 134 so what is the inequality that we get ?

OpenStudy (anonymous):

so can I do 134/2? to find L?

OpenStudy (hoblos):

before that, we must write the inequality

OpenStudy (mathmale):

What does "greater than or equal to 134" mean? Explain.

OpenStudy (anonymous):

it has to be atlease 134 or 135+

OpenStudy (mathmale):

It has to be at least 134. What has to be at least 134?

OpenStudy (anonymous):

the Total Deck

OpenStudy (mathmale):

What is the symbolic way of writing "at least 134"?

OpenStudy (mathmale):

We're discussing the dimensions of the deck, remember? Not "the total deck."

OpenStudy (anonymous):

</=

OpenStudy (mathmale):

From earlier: 58+2L represents the PERIMETER of the deck. The perimeter must be at least 134. Write the appropriate inequality.

OpenStudy (hoblos):

"<" is LESS than ">" is GREATER than

OpenStudy (anonymous):

hmm....so 58+76?

OpenStudy (hoblos):

we are trying to find an inequality we find that P = 58 + 2L and we know that P >= 134 if we replace P by 58 + 2L in the inequality, what do we get ?

OpenStudy (anonymous):

so 134>=58+2L?

OpenStudy (hoblos):

what you wrote means: "134 is greater or equal to the perimeter" we actually need the opposite 58+2L mus be greater than or equal 134

OpenStudy (anonymous):

im confused..

OpenStudy (hoblos):

lets try again P = 58 + 2L do you understand how we got this ?

OpenStudy (anonymous):

yes

OpenStudy (hoblos):

good, the question is saying "The perimeter of the deck is to be at least 134 feet" this means P >= 134 you understand this one ?

OpenStudy (anonymous):

yes P is at lease 134 and greater then 134 correct?

OpenStudy (hoblos):

yeah, "at least 134" means "greater than or equal 134"

OpenStudy (anonymous):

ok I get it

OpenStudy (hoblos):

great so if we replace P by "58 + 2L" what will we get ?

OpenStudy (anonymous):

78

OpenStudy (hoblos):

keep it as inequality don't solve it

OpenStudy (anonymous):

58+2L>=134

OpenStudy (hoblos):

Greatt! that's the inequality that we need now if we subtract 58 from both sides, what do we get?

OpenStudy (anonymous):

2L=78

OpenStudy (hoblos):

recalculate 134-58= ?

OpenStudy (anonymous):

2L=76

OpenStudy (hoblos):

where is the ">" sign ?

OpenStudy (anonymous):

2L>=76

OpenStudy (hoblos):

Great now divide both sides by 2

OpenStudy (anonymous):

L=38

OpenStudy (mathmale):

because we're looking for an inequality in L (not an equation)

OpenStudy (hoblos):

you missed the ">" again

OpenStudy (anonymous):

gahh L>=38

OpenStudy (hoblos):

Great job!

OpenStudy (anonymous):

sweet!! thanks Hoblos! and Thanks mathmale!!

OpenStudy (hoblos):

glad to help! :)